OpenStudy (jade):

Choose the correct product of (5x + 4)2.

OpenStudy (steve816):

Do you know the distributive property?

OpenStudy (will.h):

This isn't about destrubtive

OpenStudy (jade):

no im really bad with stuff

OpenStudy (will.h):

This is the expression (5x + 4)^2 correct?

OpenStudy (jade):

yes

OpenStudy (steve816):

Ohh that makes more sense.

OpenStudy (steve816):



OpenStudy (jade):

ok

OpenStudy (will.h):

Okay that means you can also have it this form (5x + 4) (5x + 4) Now you can use destrubtive property 5x * 5x = 25x^2 5x * 4 = 20x 4 * 5x = 20 x 4 * 4 = 16 Therefore we have 25x^2 + 20x + 20x + 16 Combine alike terms 25x^2 + 40x + 16

OpenStudy (jade):

i see kinda

OpenStudy (will.h):

Good. That's all

OpenStudy (jade):

so the answer would be c right

OpenStudy (jade):

negtive 6

OpenStudy (steve816):

You never gave us choices

OpenStudy (will.h):

You didn't post options

OpenStudy (will.h):

And there is no negative sign in the answer anyway

OpenStudy (jade):

i cant post option

OpenStudy (will.h):

You don't have to cuz I already gave you the steps along with the answer

OpenStudy (steve816):

Listen to Will, he already gave you the answer lol

OpenStudy (will.h):

Lol.

OpenStudy (jade):

so this is the answer 25x^2 + 40x + 16

OpenStudy (will.h):

Yeah

OpenStudy (jade):

thank u so much

OpenStudy (will.h):

You welcome